OpenStudy (anonymous):

Please Help me!!!! Please!!!!!!!!!!!!!!! A chemistry teacher asks his students to make a 5% sodium solution. The students are given 9 liters of 15% sodium solution that they are to completely mix with water. How many liters of water must be mixed to get the 5% solution??? Does anyone know the formula??

OpenStudy (anonymous):

Please help in any way possible

OpenStudy (anonymous):

(c)(a)+(c)(a)=(combined concentrate)(total amount) ???????????????

OpenStudy (anonymous):

First, you need to find the quantity of sodium. 9liters are 9000 milliliters. So 1% of water is 9000÷100=90milliliters. You have 15% of sodium. Then 90×15=1350milliliters. When you mixed some water, 1350milliliters of sodium should be 5%. So you have to find the value of 1% of When you mixed some water. Then 1% of the value is 1350÷5=270milliliters. All of the quantity is 100%, so you should do 1%×100. Then 270×100=27000milliliters. Then you want how much increased, you do 27000−1250=25750. The answer is 25.75liters.

OpenStudy (anonymous):

26?

OpenStudy (anonymous):

yeah